# Disentangling Latent Shifts of In-Context Learning Through Self-Training

### v1

@@ Abstract
In-context learning (ICL) has become essential in natural language processing, particularly with autoregressive large language models capable of learning from demonstrations provided within the prompt. However, ICL faces challenges with stability and long contexts, especially as the number of demonstrations grows, leading to poor generalization and inefficient inference. To address these issues, we introduce sticl (Self-Training ICL), an approach that disentangles the latent shifts of demonstrations from the latent shift of the query through self-training. sticl employs a teacher model to generate pseudo-labels and trains a student model using these labels, encoded in an adapter module. The student model exhibits weak-to-strong generalization, progressively refining its predictions over time. Our empirical results show that sticl improves generalization and stability, consistently outperforming traditional ICL methods and other disentangling strategies across both in-domain and out-of-domain data.
@@ 1 Introduction
In-context learning (ICL) (Brown et al., 2020) has emerged as a significant machine learning paradigm, particularly in natural language processing (NLP) applications that utilize large language models (LLMs). Unlike traditional supervised machine learning methods that rely on training over multiple epochs with large datasets, ICL leverages the ability of autoregressive LLMs to learn from context, with demonstrations and the query combined in a single prompt. This enables models to rapidly adjust to new tasks or varying input patterns without the need for additional fine-tuning. Moreover, ICL proves effective in low-resource setups by utilizing zero-shot and few-shot learning to perform tasks with minimal or no supervision (Dong et al., 2024a).
Despite its strengths, ICL faces several critical challenges. One of the key issues is stability – autoregressive LLMs based on the transformer architecture (Vaswani et al., 2017) can be highly sensitive to variations in the input context, such as the selection and ordering of demonstrations (Li et al., 2024; Lu et al., 2021; Dong et al., 2024a). This instability can result in poor generalization, making the models less reliable in real-world applications. Compounding this issue, ICL often involves long contexts because it requires incorporating multiple demonstrations alongside the query within a single input prompt. As more demonstrations are added, the input lengthens, and LLMs often struggle to handle extended contexts effectively. This problem can be traced to inherent primacy and recency biases, which lead models to overemphasize information positioned at the beginning or end of the context (Liu et al., 2024). Moreover, the inherent limitations of the context window size impose computational constraints, presenting a practical bottleneck (Dong et al., 2024b). Even with expanded context windows in newer models, limited context remains a significant challenge for ICL. LLMs struggle to fully leverage contexts when incorporating multiple demonstrations, often exceeding practical input lengths.
The aforementioned stability issues in ICL stem from the joint processing of demonstrations and the query. Since ICL can be viewed as introducing shifts in the model’s internal representations – where knowledge from demonstrations is superimposed onto the latent features induced by the query – a promising solution is to disentangle these latent shifts, separating those induced by demonstrations from those of the query. By separating these shifts, ICL can process queries independently of demonstrations, reducing computational overhead and improving stability. Disentangling has been explored from various perspectives: Liu et al. (2023) and Zhang et al. (2024) have focused on improving ICL’s stability and scalability, while Dai et al. (2023) and Todd et al. (2024) leveraged disentangling to gain theoretical insights. Separating the latent shifts makes it possible to persistently store the context knowledge provided by demonstrations, eliminating the need to reprocess demonstrations with every query. This results in significantly shorter prompts, as only the queries remain, which can mitigate the problem of long context and improve the efficiency of inference. The latent shift induced by demonstrations can then be applied trivially, for example, by adding it to the latent features induced by the query. While disentangling the latent shifts of ICL has shown potential in improving ICL and advancing theoretical understanding, current methods rely on approximations, primarily by manipulating attention heads or hidden states. A more direct and principled approach to disentangling these shifts remains an open and compelling area for further investigation.
In this work, we propose to disentangle the latent shift of demonstrations from that of the query by explicitly focusing on the model’s final outputs through the use of self-training (Amini et al., 2022). Self-training involves training a model using pseudo-labels generated by a previously learned model and has proven highly effective in leveraging unlabeled data for neural network training (Wei et al., 2021). We employ self-training in a simple teacher-student framework to encode the latent shift of demonstrations into a small set of additional parameters housed within an adapter module (Houlsby et al., 2019). Our method, sticl (Self-Training ICL), illustrated in Figure 1, employs a teacher LLM to generate pseudo-labels by processing both the demonstrations and the query without requiring extra labeled data. These pseudo-labels are then used to train a student LLM. The student model is trained to match the output provided by the teacher, taking only the query as the input. By leveraging unlabeled data through self-training, the student can correct the pseudo-labels provided by the teacher, exhibiting weak-to-strong generalization (Lang et al., 2024). The method encodes the information from the demonstrations into the parameters and can seamlessly apply the latent shift just by activating the adapter module. Furthermore, due to the flexibility of adapters, a large set of demonstrations can be chunked into more manageable subsets, with each subset encoded in its own adapter module, and the modules can be easily merged. We evaluate sticl using autoregressive LLMs such as Llama 3 (8B) (Dubey et al., 2024) and Phi 3 (mini 4k) (Abdin et al., 2024) on the GLUE (Wang et al., 2018) and MMLU (Hendrycks et al., 2021) benchmarks, comparing it to pattern-based fine-tuning (Schick & Schütze, 2021) and few-shot ICL. On both in-domain (ID) and out-of-domain (OOD) data, sticl consistently outperforms these baselines and other disentanglement methods that leverage attention heads or hidden states, thus offering a reliable alternative without needing extra labeled data.
Our contribution is two-fold: (1) We introduce sticl, a self-training ICL method that enhances efficiency and addresses stability and long-context challenges of ICL by disentangling the latent shifts between demonstrations and queries using one or several adapter modules; (2) We empirically demonstrate that sticl significantly improves both stability and generalization on ID and OOD, outperforming traditional ICL methods and other disentangling methods, while maintaining parameter efficiency. These findings suggest that even simple self-training setups, when properly designed, can offer substantial gains in ICL performance, paving the way for more efficient and scalable alternatives to current approaches. [^1] The code is included in the supplementary material and will be made available upon publication.
Teacher LLMStudent LLMBase LLMAdapterDemonstrations:$\mathbf{X}_{d}=[\mathbf{x}_{1},\,\mathbf{x}_{2},\,\ldots,\,\mathbf{x}_{n}]$Query: $\mathbf{x}_{q}$
@@ 2.1 Disentangling Latent Shifts
Disentangling in-context knowledge from the query can aid in improving the efficiency and stability of ICL. Current approaches rely on manipulating the outputs of attention heads or hidden states. The motivation behind disentangling lies in previous research (Aizerman, 1964; Irie et al., 2022), demonstrating that linear layers optimized through gradient descent have a dual form of linear attention. To illustrate, consider a neural network’s linear layer, where $\mathbf{W}_{0},\Delta\mathbf{W}\in\mathbb{R}^{m\times n}$ denote the initial weight matrix and its subsequent updates by backpropagation, respectively. With $\mathbf{x}\in\mathbb{R}^{m}$ as the input representation, a linear transformation $\mathbf{f}:\mathbb{R}^{m}\to\mathbb{R}^{n}$ can be expressed as:
$\mathbf{f}(\mathbf{x})=(\mathbf{W}_{0}+\Delta\mathbf{W})\mathbf{x}\text{.}$ (1)
During backpropagation, $\Delta\mathbf{W}$ is computed by accumulating the outer products (denoted by $\otimes$) of $N$ training examples $\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{N}\}$, where $\mathbf{x}_{i}\in\mathbb{R}^{m}$, and the error signals $\{\mathbf{e}_{1},\mathbf{e}_{2},\dots,\mathbf{e}_{N}\}$, where $\mathbf{e}_{i}\in\mathbb{R}^{n}$, obtained from the gradients of the loss function:
$\Delta\mathbf{W}=\sum_{i=1}^{N}{\mathbf{e}_{i}\otimes\mathbf{x}_{i}}\text{.}$ (2)
Irie et al. (2022) show that the update part of linear layers optimized by gradient descent can be expressed as unnormalized linear dot-product attention:
[…]
In the context of the attention mechanism, this shows that the latent shift $\Delta\mathbf{W}\mathbf{x}$ corresponds directly to the application of linear attention, with error signals $\mathbf{e}_{i}$ as values, training examples $\mathbf{x}_{i}$ as keys, and the current input $\mathbf{x}$ as the attention query.
The concept of disentangling the latent shifts described in (3) can be extended to ICL, albeit only under the approximation of linear attention. Let $\mathbf{W}_{V}$, $\mathbf{W}_{K}$, and $\mathbf{W}_{Q}$ denote the weight matrices for values, keys, and queries, respectively. Let $\mathbf{x}_{q}^{(t)}$ represent the current query token’s embedding at step $t$, and $\mathbf{q}^{(t)}=\mathbf{W}_{Q}\mathbf{x}_{q}^{(t)}$ is the corresponding attention query vector. The matrix $\mathbf{X}_{q}=[\mathbf{x}_{q}^{(1)},\mathbf{x}_{q}^{(2)},\dots,\mathbf{x}_{q}^{(t-1)}]$ contains all previous query token representations up to $t-1$, and $\mathbf{X}_{d}$ is the matrix of demonstration token representations. The concatenation $[\mathbf{X}_{d};\mathbf{X}_{q}]$ along the sequence dimension is used to compute the attention output at step $t$, expressed as:
@@
where $d$ is the scaling factor (i.e., the dimensionality of the key vectors). By approximating the attention mechanism with linear attention, it becomes possible to disentangle the latent shift of the zero-shot output of an attention head induced by the query from the latent shift induced by the demonstrations (Dai et al., 2023):
@@
$=\underbrace{\mathbf{W}_{V}\mathbf{X}_{q}\left(\mathbf{W}_{K}\mathbf{X}_{q}\right)^{\top}}_{\mathbf{W}_{\text{ZS}}}\mathbf{q}^{(t)}+\underbrace{\mathbf{W}_{V}\mathbf{X}_{d}\left(\mathbf{W}_{K}\mathbf{X}_{d}\right)^{\top}}_{\Delta\mathbf{W}_{\text{ICL}}}\mathbf{q}^{(t)}$
$=\left(\mathbf{W}_{\text{ZS}}+\Delta\mathbf{W}_{\text{ICL}}\right)\mathbf{q}^{(t)}\text{.}$
@@
This approximation disentangles the latent shift induced by the demonstrations $\mathbf{X}_{d}$ from that induced by the query $\mathbf{x}_{q}^{(t)}$ (cf. Appendix A for detailed derivation of (5)). The contribution from ICL is captured as a virtual weight update $\Delta\mathbf{W}_{\text{ICL}}$, corresponding to virtual gradients, often referred to as “meta-gradients” in the literature. The zero-shot latent shift of the query, corresponding to $\mathbf{W}_{\text{ZS}}\mathbf{q}^{(t)}$, reflects the output without demonstrations, providing the initial state. Analogous to $\Delta\mathbf{W}\mathbf{x}$ in (3), the latent shift $\Delta\mathbf{W}_{\text{ICL}}\mathbf{q}^{(t)}$ reflects the contribution of ICL. Finally, by substituting $\mathbf{h}_{\text{ZS}}=\mathbf{W}_{\text{ZS}}\mathbf{q}^{(t)}$ and $\Delta\mathbf{h}_{\text{ICL}}=\Delta\mathbf{W}_{\text{ICL}}\mathbf{q}^{(t)}$, we can rewrite the output of an attention head as:
$\mathbf{f}_{\text{AH}}(\mathbf{x}_{q}^{(t)})\approx\mathbf{h}_{\text{ZS}}+\Delta\mathbf{h}_{\text{ICL}}\text{.}$ (6)
@@
Although transformer-based LLMs use non-linear attention in practice, many approaches (Dai et al., 2023; Zhang et al., 2024; Todd et al., 2024) rely on the theoretical underpinnings of linear attention. These methods manipulate attention heads or hidden states to disentangle latent shifts despite the inherent non-linearity of the models. Furthermore, this simplification overlooks other crucial components of the transformer architecture, such as the feed-forward layers, activation functions, and residual connections. While approaches based on linear attention have proven effective, they leave room for further improvements in capturing and disentangling the full complexity of how transformers process data. In this work, we explore how virtual weight updates can be obtained more directly while preserving the key components of the transformer architecture.
2.2 Self-Training ICL
Building on the concept of disentangling latent shifts in transformer architectures, we introduce sticl (Self-Training ICL), an approach that offers a simple yet highly efficient way to internalize ICL knowledge into a model’s parameters. Rather than relying solely on manipulating attention heads, as is common in current methods, sticl aims to capture the full complexity of the transformer’s components – considering the final output, which depends on all layers, including attention heads, feed-forward layers, and residual connections. By aligning more directly with the actual latent shifts induced by demonstrations, sticl ensures that the model uses the entirety of its architecture to first embed and later apply in-context knowledge.
@@ 2.2 Self-Training ICL
At the core of sticl is a simple teacher-student framework: the teacher model, $\mathbf{f}_{\text{teacher}}$, processes both demonstrations and the query together to generate pseudo-labels without needing additional labeled data. The student model, $\mathbf{f}_{\text{student}}$, shares the same architecture as the teacher but includes adapter parameters. Unlike the teacher, the student processes only the query, using the adapter to internalize the knowledge from the demonstrations, as illustrated in Figure 1. Let $\mathbf{x}_{q}$ denote the query input and $\mathbf{X}_{d}$ the matrix of demonstration tokens, where each row corresponds to a single demonstration. [^2] Since the embedding transformation is implied within the LLM, the query $\mathbf{x}_{q}$ is a vector of token IDs, and $\mathbf{X}_{d}$ contains token IDs of demonstrations. The empirical loss, using the cross-entropy loss $\ell_{\text{CE}}$, is defined as:
@@
where $\mathcal{D}_{\text{unlab}}$ is an unlabeled dataset and $\mathbf{X}^{*}_{d}$ is a flattened version of $\mathbf{X}_{d}$. This approach is grounded in self-training (Amini et al., 2022), leveraging the teacher’s pseudo-labels to fine-tune the student’s adapter.
sticl fundamentally differs from existing approaches, which rely on manipulating attention heads or hidden states at query time. Instead, sticl progressively embeds the knowledge from demonstrations into the adapter parameters, denoted $\mathbf{W}_{\text{ICL}}$. The base LLM parameters, $\mathbf{W}_{\text{ZS}}$, capture the zero-shot component, while the total model parameters may be represented as $\mathbf{W}_{\text{ZS}}\oplus\mathbf{W}_{\text{ICL}}$, where $\oplus$ denotes the composition of base and adapter parameters. [^3] Notably, the number of adapter parameters is significantly smaller compared to the base model parameters. This setup captures the latent shift introduced by the demonstrations through $\mathbf{W}_{\text{ICL}}$, extending the disentangling process outlined by (5) across the model’s entire architecture. The teacher processes the full input sequence $\left[\mathbf{X}^{*}_{d};\mathbf{x}_{q}\right]$, while the student processes only the query, applying $\mathbf{W}_{\text{ICL}}$ to integrate demonstration knowledge without explicitly processing the demonstrations. Analogous to (6), the latent shift induced by demonstrations can be recovered by decomposing outputs into zero-shot and ICL components. Let $\mathbf{h}_{\text{LLM}}(\mathbf{x}_{q}\mid\mathbf{W})$ represent the final latent states of an LLM with parameters $\mathbf{W}$ when processing the input $\mathbf{x}_{q}$. The following decomposition holds:
@@
where $\Delta\mathbf{h}_{\text{ICL}}$ encapsulates the latent shift attributable to the demonstrations. sticl encodes the latent shift implicitly within the adapter parameters $\mathbf{W}_{\text{ICL}}$, which is central to our approach. However, if necessary, the latent shift can also be explicitly calculated owing to the decomposition in (8).
The stabilizing effect of sticl extends beyond just handling demonstrations. By iterating over multiple epochs, sticl leverages the same LLM instance for both the teacher and student roles, transitioning smoothly between them by activating or deactivating the adapter. Demonstrations can be shuffled across epochs to reduce sensitivity to their order, further stabilizing the ICL process. But the true power of sticl emerges from its parametric nature, which aligns with the optics of weak-to-strong generalization (Lang et al., 2024). The adapter parameters allow the model to internalize shifts and generalize effectively across both ID and OOD data, as demonstrated empirically in our experiments (cf. Section 3).
From the perspective of weak-to-strong generalization, the student model is not just expected to match the teacher – it is designed to outperform it. sticl facilitates this by leveraging pseudo-label correction, where incorrect labels are refined using high-confidence neighboring examples, and coverage expansion, enabling the model to generalize beyond regions initially covered by the teacher and even to near-OOD data (Section 3). sticl not only stabilizes ICL but also capitalizes on the parametric regime, where latent shifts can be efficiently encoded, enabling the model to establish implicit local-to-global consistency across the data distribution through extrapolation (Wei et al., 2021).

### v2

# Disentangling Latent Shifts of In-Context Learning with Weak Supervision

@@ Abstract
In-context learning (ICL) enables large language models to perform few-shot learning by conditioning on labeled examples in the prompt. Despite its flexibility, ICL suffers from instability – especially as prompt length increases with more demonstrations. To address this, we treat ICL as a source of weak supervision and propose a parameter-efficient method that disentangles demonstration-induced latent shifts from those of the query. An ICL-based teacher generates pseudo-labels on unlabeled queries, while a student predicts them using only the query input, updating a lightweight adapter. This captures demonstration effects in a compact, reusable form, enabling efficient inference while remaining composable with new demonstrations. Although trained on noisy teacher outputs, the student often outperforms its teacher through pseudo-label correction and coverage expansion, consistent with the weak-to-strong generalization effect. Empirically, our method improves generalization, stability, and efficiency across both in-domain and out-of-domain tasks, surpassing standard ICL and prior disentanglement methods.
@@ 1 Introduction
In-context learning (ICL) has become a core mechanism for adapting large language models (LLMs) to new tasks in a way that removes the need to update their parameters [3, 9]. By prepending a few labeled examples, called demonstrations, to the input query, LLMs can perform few-shot learning directly at inference time. This paradigm is especially attractive in low-resource settings, where full fine-tuning is too costly or impractical.
Despite its convenience, ICL performance is highly sensitive to the selection and ordering of demonstrations, often resulting in unstable predictions and poor generalization [29, 24]. Moreover, ICL typically requires long contexts, as multiple demonstrations must be included alongside the query in a single input. As input lengths grow, inference becomes increasingly inefficient, inflating processing costs, amplifying positional biases – including inherent primacy and recency effects in transformer-based LLMs [26] – and pushing against the model’s context window limits [10]. Consequently, ICL scales poorly with the number of demonstrations [5, 4]: beyond a certain threshold, additional examples either degrade performance or must be discarded entirely. This inefficiency and poor scalability limit ICL’s ability to incorporate more supervision, preventing it from fully leveraging the potential benefits of richer demonstrations.
To address these limitations, adopting a mechanistic perspective on ICL has proven useful. In this view, demonstrations influence model behavior by inducing latent shifts – context-dependent changes in internal representations that alter how the model processes the query. By disentangling these shifts from the representation of the query itself, ICL can operate more robustly, processing queries independently of demonstrations. This, in turn, enables contextual knowledge to be stored persistently, eliminating the need to reprocess demonstrations for each new query. The latent shifts can then be reapplied directly, thereby reducing prompt length, improving inference efficiency, and enabling more modular and reusable representations. The disentanglement of demonstration-induced latent shifts from those of the query has been explored from both practical and theoretical perspectives: some work aims to improve the stability and scalability of ICL [27, 41], while other studies provide formal insights into the nature of context-induced shifts [7, 33]. However, existing methods rely on approximations by manipulating attention heads or hidden states. Developing a more direct approach to disentangling these shifts remains a compelling direction for future work.
In this paper, we propose to disentangle the latent shifts of the demonstrations and the query by using the output of ICL as a weak supervision signal, rather than approximating the underlying internal mechanisms. Specifically, we use ICL predictions as pseudo-labels that capture the full, contextualized influence of demonstrations. These pseudo-labels guide the training of a student model that internalizes demonstration-conditioned behavior – without repeated prompting or architectural intervention. We instantiate this idea with wilda (Weakly-supervised In-context Learning Disentanglement via Adapters), a parameter-efficient method for encoding the latent shifts induced by in-context demonstrations. In a teacher–student setup, an ICL-based teacher generates pseudo-labels for unlabeled queries, while the student learns to predict them using only the query input. The student updates a lightweight adapter module [15], enabling it to capture the shift in a reusable and modular form. Once trained, the adapter supports efficient inference without requiring demonstrations in the prompt. At inference time, additional demonstrations can still be provided, and are composed with the adapter’s shift – enabling flexible combinations of prompt-based and parameter-based task control.
We evaluate wilda on both in-domain (ID) and out-of-domain (OOD) data, comparing it to standard ICL, prompt-based fine-tuning [32], and recent approaches that manipulate architecture or hidden states to disentangle latent shifts. wilda consistently improves generalization, prompt robustness, and inference efficiency, while remaining highly parameter-efficient. Despite learning from ICL outputs, the student often surpasses its teacher. We experimentally show that this improvement arises from two emergent behaviors: pseudo-label correction, where the student refines noisy or inconsistent outputs from the teacher, and coverage expansion, where it generalizes beyond the narrow patterns encoded in the demonstrations. Together, these effects enable weak-to-strong (W2S) generalization [22], allowing the model to learn stable task behavior from limited supervision.
Our contribution is twofold: (1) we propose wilda, a method that encodes ICL-induced behavior into reusable adapters, improving inference efficiency and prompt stability without requiring demonstration prompts; and (2) we show that wilda outperforms traditional ICL and latent shift disentanglement methods on both ID and OOD data, while maintaining parameter efficiency. Together, these results show that treating ICL as a source of weak supervision enables robust and scalable behavior modeling – offering an effective alternative to ICL inference.
[…]
@@ 2.1 Disentangling Latent Shifts
Disentangling in-context knowledge from the query can enhance the efficiency and stability of ICL. Current methods typically achieve disentanglement by modifying the outputs of attention heads or hidden states. The theoretical motivation for this approach stems from previous studies [2, 19], which demonstrate that linear layers optimized via gradient descent can be viewed through the lens of linear attention mechanisms. Specifically, consider a neural network’s linear layer characterized by an initial weight matrix $\mathbf{W}_{0}\in\mathbb{R}^{m\times n}$ and an update $\Delta\mathbf{W}\in\mathbb{R}^{m\times n}$ resulting from backpropagation. Given an input representation $\mathbf{x}\in\mathbb{R}^{m}$, the linear transformation $\mathbf{f}:\mathbb{R}^{m}\to\mathbb{R}^{n}$ can be expressed succinctly as $\mathbf{f}(\mathbf{x})=(\mathbf{W}_{0}+\Delta\mathbf{W})\mathbf{x}$. Let $\mathbf{x}_{i}\in\mathbb{R}^{m}$ be a training example and $\mathbf{e}_{i}\in\mathbb{R}^{n}$ the error signal on $\mathbf{x}_{i}$ obtained from the gradient of the loss function. During backpropagation, $\Delta\mathbf{W}$ is computed by accumulating the outer products (denoted by $\otimes$) of $N$ training examples $\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{N}\}$ and their error signals $\{\mathbf{e}_{1},\mathbf{e}_{2},\dots,\mathbf{e}_{N}\}$, i.e., $\Delta\mathbf{W}=\sum_{i=1}^{N}\mathbf{e}_{i}\otimes\mathbf{x}_{i}$. The update part of linear layers optimized by gradient descent can be expressed as unnormalized linear dot-product attention [19]:
$\mathbf{f}(\mathbf{x})=(\mathbf{W}_{0}+\Delta\mathbf{W})\mathbf{x}=\mathbf{W}_{0}\mathbf{x}+\sum_{i=1}^{N}{(\mathbf{e}_{i}\otimes\mathbf{x}_{i})\mathbf{x}}=\mathbf{W}_{0}\mathbf{x}+\underbrace{\sum_{i=1}^{N}{\mathbf{e}_{i}(\mathbf{x}_{i}^{T}\mathbf{x}})}_{\text{linear attention}}.$ (1)
In the context of the attention mechanism, this shows that the latent shift $\Delta\mathbf{W}\mathbf{x}$ induced by training examples corresponds directly to the application of linear attention, with error signals $\mathbf{e}_{i}$ as values, training examples $\mathbf{x}_{i}$ as keys, and the current input $\mathbf{x}$ as the attention query.
The concept of disentangling the latent shifts described in (1) can be extended to ICL, albeit only under the approximation of linear attention. Let $\mathbf{W}_{V}$, $\mathbf{W}_{K}$, and $\mathbf{W}_{Q}$ denote the weight matrices for values, keys, and queries, respectively. Let $\mathbf{x}_{q}^{(t)}$ represent the current query token’s embedding at step $t$, and $\mathbf{q}^{(t)}=\mathbf{W}_{Q}\mathbf{x}_{q}^{(t)}$ is the corresponding attention query vector. The matrix $\mathbf{X}_{q}=[\mathbf{x}_{q}^{(1)},\mathbf{x}_{q}^{(2)},\dots,\mathbf{x}_{q}^{(t-1)}]$ contains all previous query token representations up to t1t-1, and $\mathbf{X}_{d}$ is the matrix of demonstration token representations. The concatenation $\mathbf{X}_{d};\mathbf{X}_{q}]$ along the sequence dimension is used to compute the output of a single attention head (AH) at step $t$, expressed as:
@@
where $d$ is the scaling factor (i.e., the dimensionality of the key vectors). By approximating the attention mechanism with linear attention, it becomes possible to disentangle the latent shift of the zero-shot output of an attention head induced by the query from the latent shift induced by the demonstrations [7]:
@@
$=\underbrace{\mathbf{W}_{V}\mathbf{X}_{q}\left(\mathbf{W}_{K}\mathbf{X}_{q}\right)^{\top}}_{\mathbf{W}_{\text{ZS}}}\mathbf{q}^{(t)}+\underbrace{\mathbf{W}_{V}\mathbf{X}_{d}\left(\mathbf{W}_{K}\mathbf{X}_{d}\right)^{\top}}_{\Delta\mathbf{W}_{\text{ICL}}}\mathbf{q}^{(t)}.$
@@
This approximation disentangles the latent shift induced by the demonstrations $\mathbf{X}_{d}$ from that induced by the query $\mathbf{x}_{q}^{(t)}$ (see Appendix A for a detailed derivation of (3)). The contribution from ICL is captured as a virtual weight update $\Delta\mathbf{W}_{\text{ICL}}$, corresponding to virtual gradients, often referred to as “meta-gradients” in the literature. The zero-shot latent shift of the query, corresponding to $\mathbf{W}_{\text{ZS}}\mathbf{q}^{(t)}$, reflects the output without demonstrations, providing the initial state. Analogous to $\Delta\mathbf{W}\mathbf{x}$ in (1), the latent shift $\Delta\mathbf{W}_{\text{ICL}}\mathbf{q}^{(t)}$ reflects the contribution of ICL. Finally, by substituting $\mathbf{h}_{\text{ZS}}=\mathbf{W}_{\text{ZS}}\mathbf{q}^{(t)}$ and $\Delta\mathbf{h}_{\text{ICL}}=\Delta\mathbf{W}_{\text{ICL}}\mathbf{q}^{(t)}$, we can rewrite the output of an attention head as:
$\mathbf{f}_{\text{AH}}(\mathbf{x}_{q}^{(t)})\approx\mathbf{h}_{\text{ZS}}+\Delta\mathbf{h}_{\text{ICL}}.$ (4)
@@
Although transformer-based LLMs employ non-linear attention in practice, many methods [7, 41, 33] rely on theoretical assumptions from linear attention, manipulating attention heads or hidden states to approximate latent shift disentanglement. This simplification, however, overlooks key architectural components such as feed-forward layers, activation functions, and residual connections. While effective to a degree, these methods fall short of fully capturing the complex dynamics through which transformers process demonstrations. In this work, we explore how virtual weight updates can be obtained more directly while preserving the key components of the transformer architecture.
2.2 Weak Supervision with ICL
To disentangle the latent shifts induced by in-context demonstrations, we introduce wilda, a method that uses ICL predictions as a form of weak supervision to encode these shifts into reusable adapter parameters [15, 16]. Instead of focusing narrowly on attention head manipulations, wilda captures the full impact of demonstrations as expressed in the model’s final outputs – reflecting the combined effects of all components, including attention layers, feed-forward blocks, and residual paths. By aligning with the actual latent shifts induced by ICL, wilda enables the model to embed and reapply in-context knowledge using its full architecture, without relying on repeated prompting.
@@ 2.2 Weak Supervision with ICL
At the core of wilda is a simple teacher–student framework: the teacher model, $\mathbf{f}_{\text{teacher}}$, processes both the demonstrations and the query together to generate pseudo-labels without requiring additional labeled data. The student model, $\mathbf{f}_{\text{student}}$, shares the same architecture as the teacher but includes adapter parameters. Unlike the teacher, the student processes only the query, using the adapter to internalize the knowledge from the demonstrations, as illustrated in Figure 1. Let $\mathbf{x}_{q}$ denote the query input and $\mathbf{X}_{d}$ the matrix of demonstration tokens, where each row corresponds to a single demonstration. The empirical loss is defined using the cross-entropy loss function $\ell_{\text{CE}}$, which aligns the student’s output distribution with the teacher’s full probability distribution over the vocabulary. This enables the student to learn from the full signal provided by the teacher’s output logits. Formally, the empirical loss is
@@
where $\mathcal{D}_{\text{unlab}}$ is an unlabeled dataset and $\mathbf{X}^{*}_{d}$ is a flattened version of $\mathbf{X}_{d}$.
wilda fundamentally differs from existing approaches, which manipulate attention heads or hidden states at query time, by instead progressively embedding the knowledge from demonstrations into the adapter parameters, denoted $\mathbf{W}_{\text{ICL}}$. The base LLM parameters, $\mathbf{W}_{\text{ZS}}$, capture the zero-shot component, while the total model parameters may be represented as $\mathbf{W}_{\text{ZS}}\oplus\mathbf{W}_{\text{ICL}}$, where $\oplus$ denotes the composition of base and adapter parameters. [^1] Notably, the number of adapter parameters is significantly smaller compared to the base model parameters. This setup captures the latent shift introduced by the demonstrations through $\mathbf{W}_{\text{ICL}}$, extending the disentangling process outlined by (3) across the model’s entire architecture. The teacher processes the full input sequence $\left[\mathbf{X}^{*}_{d};\mathbf{x}_{q}\right]$, while the student processes only the query, applying $\mathbf{W}_{\text{ICL}}$ to integrate demonstration knowledge without explicitly processing the demonstrations. Analogously to (4), the latent shift induced by demonstrations can be recovered by decomposing outputs into zero-shot and ICL components. Let $\mathbf{h}_{\text{LLM}}(\mathbf{x}_{q}\mid\mathbf{W})$ represent the final latent states of an LLM with parameters $\mathbf{W}$ when processing the input $\mathbf{x}_{q}$. The following decomposition holds:
@@
where $\Delta\mathbf{h}_{\text{ICL}}$ encapsulates the latent shift attributable to the demonstrations. wilda encodes the latent shift implicitly within the adapter parameters $\mathbf{W}_{\text{ICL}}$, which is central to our approach. However, if necessary, the latent shift can also be explicitly calculated owing to the decomposition in (6).
wilda achieves stability not only by disentangling demonstration effects, but also through its training dynamics and parametric nature. During training, the same LLM instance serves as both teacher and student across epochs, with the adapter toggled on or off to alternate between roles. Shuffling demonstrations between epochs mitigates order sensitivity, further stabilizing the ICL process. Crucially, wilda leverages its parametric adapter to internalize demonstration-induced shifts, enabling the model to generalize effectively across ID and near-OOD data (see Section 3). This aligns naturally with the W2S generalization paradigm [22], in which the student is not merely expected to match the teacher but to surpass it. wilda facilitates this process by compactly encoding latent shifts in a way that supports both pseudo-label correction (refining noisy targets through local consistency) and coverage expansion (generalizing beyond the teacher’s original scope). Together, these effects enable stable extrapolation across the data distribution, aligning with theoretical expectations of W2S generalization [38].